OpenStudy (anonymous):

MEDAL The sum of 1/9, 2/3, and 5/18is A. 12/9. B. 4/15. C. 19/18. D. 8/30.

OpenStudy (anonymous):

@ZambieMan

OpenStudy (studygurl14):

\[\frac{ 1 }{ 9 } + \frac{ 2 }{ 3 } + \frac{ 5 }{ 18 }\] \[(\frac{ 1 }{ 9})\frac{ 2 }{ 2 } + (\frac{ 2 }{ 3 })\frac{ 3 }{ 3 } + \frac{ 5 }{ 18 }\]

OpenStudy (studygurl14):

\[\frac{ 2 }{ 18 } + \frac{ 6 }{ 18 } + \frac{ 5 }{ 18 }\]

OpenStudy (pratyush5):

@StudyGurl14 has made all the denominators common, ie 18

OpenStudy (anonymous):

\[\frac{ 13 }{ 18 }\]

OpenStudy (pratyush5):

This way we can add all the numerators.

OpenStudy (anonymous):

is that correct?

OpenStudy (pratyush5):

Yup :)

OpenStudy (anonymous):

hurrah!!

OpenStudy (pratyush5):

Hey Hey wait

OpenStudy (pratyush5):

Its wrong

OpenStudy (pratyush5):

Sorry to crash the party

OpenStudy (anonymous):

it is?

OpenStudy (anonymous):

that wasn't even an answer choice ha

OpenStudy (pratyush5):